What are the points on a graph for 14x-15y=-1

Found 2 solutions by richwmiller, math_helper:
Answer by richwmiller(17219)

Pick some x and solve for y
Pick some y and solve for x
Good choices usually include 0 for x and then 0 for y

Answer by math_helper(2461)

What are the points on a graph for 14x-15y=-1
===================================
The question should be revised. There are an infinite number of points on the line 14x-15y=-1.
-15y = -14x - 1
y = (14/15)x + (1/15)
Slope is 14/15
y-intercept is at 1/15 (i.e. when x = 0, y = 1/15)
x-intercept is at -1/14

Graph: